# Node Duplication Improves Cold-start Link Prediction

### v1

@@ 3.1 Proposed Method
Figure 2: Ablation study of NodeDup on Physics. Step (2) and Step (3) are the steps introduced in Section 3.1. Both steps play an important role in performance improvements of NodeDup.
@@
3.2 How does node duplication help cold-start LP?
In this subsection, we analyze how such a simple method can improve cold-start LP from two perspectives: the neighborhood aggregation in GNNs and the supervision signal during training. In short, NodeDup leverages the extra information from an additional “view”. The existing view is when a node is regarded as the anchor node during message passing, whereas the additional view is when that node is regarded as one of its neighbors thanks to the duplicated node from NodeDup.
Aggregation. As described in Section 2.2, when UPDATE()($\cdot$)( ⋅ ) and AGG()($\cdot$)( ⋅ ) do not share the transformation for node features, GNNs would have separate weights for self-representation and neighbor representations. The separate weights enable the neighbors and the node itself to play distinct roles in the UPDATE step. By leveraging this property, with NodeDup, the model can leverage the two “views” for each node: first, the existing view is when a node is regarded as the anchor node during message passing, and the additional view is when that node is regarded as one of its neighbors thanks to the duplicated node from NodeDup. Taking the official PyG (Fey & Lenssen, 2019) implementation of GSage (Hamilton et al., 2017) as an example, it updates node representations using $\bm{h}_{v}^{(l+1)}=\text{W}_{1}\bm{h}_{v}^{(l)}+\text{W}_{2}\bm{m}_{v}^{(l)}$. Here, $\text{W}_{1}$ and $\text{W}_{2}$ correspond to the self-representation and neighbors’ representations, respectively. Without NodeDup, isolated nodes $\mathcal{V}_{iso}$ have no neighbors, which results with $\bm{m}_{v}^{(l)}=\bm{0}$. Thus, the representations of all $v\in\mathcal{V}_{iso}$ are only updated by $\bm{h}_{v}^{(l+1)}=\text{W}_{1}\bm{h}_{v}^{(l)}$. With NodeDup, the updating process for isolated node $v$ becomes $\bm{h}_{v}^{(l+1)}=\text{W}_{1}\bm{h}_{v}^{(l)}+\text{W}_{2}\bm{h}_{v}^{(l)}=(\text{W}_{1}+\text{W}_{2})\bm{h}_{v}^{(l)}$. It indicates that $\text{W}_{2}$ is also incorporated into the node updating process for isolated nodes, which offers an additional perspective for isolated nodes’ representation learning. Similarly, GAT (Veličković et al., 2017) updates node representations with $\bm{h}_{v}^{(l+1)}=\alpha_{vv}\Theta\bm{h}^{(l)}_{v}+\sum_{u\in\mathcal{N}_{v}}\alpha_{vu}\Theta\bm{h}^{(l)}_{u}$, where $\alpha_{vu}=\frac{\text{exp}(\text{LeakyReLU}(\bm{a}^{\top}[\Theta\bm{h}^{(l)}_{v}||\Theta\bm{h}^{(l)}_{u}]))}{\sum_{i\in\mathcal{N}_{v}\cup v}\text{exp}(\text{LeakyReLU}(\bm{a}^{\top}[\Theta\bm{h}^{(l)}_{v}||\Theta\bm{h}^{(l)}_{i}]))}$. Attention scores in $\bm{a}$ partially correspond to the self-representation $\bm{h}_{v}$ and partially to neighbors’ representation $\bm{h}_{u}$. In this case, neighbor information also offers a different perspective compared to self-representation. Such “multi-view” enriches the representations learned for the isolated nodes in a similar way to how ensemble methods work (Allen-Zhu & Li, 2020). Moreover, apart from addressing isolated nodes, the same mechanism and multi-view perspective also apply to Low-degree nodes.
Supervision. For LP tasks, besides the aggregation, edges also serve as supervised training signals. Cold nodes have few or no positive training edges connecting to them, potentially leading to out-of-distribution (OOD) issues (Wu et al., 2022), especially for isolated nodes. The additional edges, added by NodeDup to connect cold nodes with their duplicates, serve as additional positive supervision signals for LP. More supervision signals for cold nodes usually lead to better-quality embeddings.
Ablation Study. Figure 2 shows an ablation study on these two designs where NodeDup w/o Step (3) indicates only using the augmented nodes and edges in aggregation but not supervision; NodeDup w/o Step (2) indicates only using the augmented edges in supervision but not aggregation. We can observe that using augmented nodes/edges either in supervision or aggregation can significantly improve the LP performance on Isolated nodes, and NodeDup, by combining them, results in larger improvements. Besides, NodeDup also achieves improvements on Low-degree nodes while not sacrificing Warm nodes’ performance.
3.3 Relation between NodeDup and Self-distillation
@@
Recently, Allen-Zhu & Li (2020) showed that the success of self-distillation, similar to our method, contributes to ensemble learning by providing models with different perspectives on the knowledge. Building on this insight, we show an interesting interpretation of NodeDup, as a simplified and enhanced version of self-distillation for LP tasks for cold nodes, illustrated in Figure 3, in which we draw a connection between self-distillation and NodeDup.
@@
Figure 4: Performance with different training strategies introduced in Figure 3 on Citeseer. NodeDup achieves better performance across all settings.
@@
In self-distillation, a teacher GNN is first trained to learn the node representations $\mathbf{H}^{t}$ from original features $\mathbf{X}$. We denote the original graph as $G^{o}$, and we denote the graph, where we replace the node features in $G^{o}$ with $\mathbf{H}^{t}$, as $G^{t}$ in Figure 3. The student GNN is then initialized with random parameters and trained with the sum of two loss functions: $\mathcal{L}_{SD}=\mathcal{L}_{SP}+\mathcal{L}_{KD}$, where $\mathcal{L}_{SP}$ denotes the supervised training loss with $G^{o}$ and $\mathcal{L}_{KD}$ denotes the knowledge distillation loss with $G^{t}$. Figure 4 shows that self-distillation outperforms the teacher GNN across all settings.
@@
The effect of $\mathcal{L}_{KD}$ is similar to that of creating an additional link connecting nodes in $G^{o}$ to their corresponding nodes in $G^{t}$ when optimizing with $\mathcal{L}_{SP}$. This is illustrated by the red dashed line in Figure 3. For better clarity, we show the similarities between these two when we use the inner product as the decoder for LP with the following example. Given a node $v$ with normalized teacher embedding $\bm{h}_{v}^{t}$ and normalized student embedding $\bm{h}_{v}$, the additional loss term that would be added for distillation with cosine similarity is $\mathcal{L}_{KD}=-\frac{1}{N}\sum_{v\in\mathcal{V}}\bm{h}_{v}\cdot\bm{h}_{v}^{t}$. On the other hand, for the dashed line edges in Figure 3, we add an edge between the node $v$ and its corresponding node $v^{\prime}$ in $G^{t}$ with embedding $\bm{h}_{v^{\prime}}^{t}$. When trained with an inner product decoder and binary cross-entropy loss, it results in the following: $\mathcal{L}_{SP}=-\frac{1}{N}\sum y_{vv^{\prime}}\log(\bm{h}_{v}\cdot\bm{h}_{v^{\prime}}^{t})+(1-y_{vv^{\prime}})\log(1-\bm{h}_{v}\cdot\bm{h}_{v^{\prime}}^{t})$. Since we always add the edge $(v,v^{\prime})$, we know $y_{vv^{\prime}}=1$, and can simplify the loss as follows: $\mathcal{L}_{SP}=-\frac{1}{N}\sum\log(\bm{h}_{v}\cdot\bm{h}_{v^{\prime}}^{t})$. Here, we can observe that $\mathcal{L}_{KD}$ and $\mathcal{L}_{SP}$ are positively correlated as $\log(\cdot)$ is a monotonically increasing function.
@@
To further improve this step and mitigate potential noise in $G^{t}$, we explore a whole graph duplication technique, where $G^{t}$ is replaced with an exact duplicate of $G^{o}$ to train the student GNN. The results in Figure 4 demonstrate significant performance enhancement achieved by whole graph duplication compared to self-distillation. NodeDup is a lightweight variation of the whole graph duplication technique, which focuses on duplicating only the cold nodes and adding edges connecting them to their duplicates. From the results, it is evident that NodeDup consistently outperforms the teacher GNN and self-distillation in all scenarios. Additionally, NodeDup exhibits superior performance on isolated nodes and is much more efficient compared to the whole graph duplication approach.
@@ 3.4 NodeDup(L): An Efficient Variant of NodeDup
Inspired by the above analysis, we further introduce a lightweight variant of NodeDup for better efficiency, NodeDup(L). To provide above-described “multi-view” information as well as the supervision signals for cold nodes, NodeDup(L) simply add additional self-loop edges for the cold nodes into the edge set $\mathcal{E}$, that is, =$\mathcal{E}^{\prime}=\mathcal{E}\cup\{e_{vv}:\forall v\in\mathcal{V}_{cold}\}$. NodeDup(L) preserves the two essential designs of NodeDup while avoiding the addition of extra nodes, which further saves time and space complexity. Moreover, NodeDup differs from NodeDup(L) since each duplicated node in NodeDup will provide another view for itself because of dropout layers, which leads to different performance as shown in Section 4.2.
NodeDup(L) vs. Self-loop. We remark upon a resemblance between NodeDup(L) and self-loops in GNNs (e.g., the additional self-connection in the normalized adjacency matrix by GCN) as they both add self-loop edges. However, they differ in two aspects. During aggregation: NodeDup(L) intentionally incorporates the self-representation $\bm{h}_{v}^{(l)}$ into the aggregated neighbors’ representation $\bm{m}^{(l)}_{v}$ by adding additional edges. Taking GSage as an example, the weight matrix $\text{W}_{2}$ would serve an extra “view” of $\bm{h}_{v}^{(l)}$ when updating $\bm{h}_{v}^{(l+1)}$, whereas the default self-loops only use information from $\text{W}_{1}$. Additionally, in the supervision signal: different from the normal self-loops and the self-loops introduced in the previous works (Cai et al., 2019; Wang et al., 2020), where self-loops are solely for aggregation, the edges added by NodeDup(L) also serve as additional positive training samples for the cold nodes.
@@ 4 Experiments
Table 1: Performance compared with base GNN model and baselines for cold-start methods(evaluated by Hits@10). The best result is bold, and the runner-up is underlined. NodeDup and NodeDup(L) outperform GSage and cold-start baselines almost all the cases.
@@
4.1 Experimental Settings
Datasets and Evaluation Settings. We conduct experiments on 7 benchmark datasets: Cora, Citeseer, CS, Physics, Computers, Photos and IGB-100K, with their details specified in Appendix B. We randomly split edges into training, validation, and testing sets. We allocated 10% for validation and 40% for testing in Computers and Photos, 5%/10% for testing in IGB-100K, and 10%/20% in other datasets. We follow the standard evaluation metrics used in the Open Graph Benchmark (Hu et al., 2020) for LP, in which we rank missing references higher than 500 negative reference candidates for each node. The negative references are randomly sampled from nodes not connected to the source node. We use Hits@10 as the main evaluation metric (Han et al., 2022) and also report MRR performance in Appendix D. We follow Guo et al. (2022b) and Shiao et al. (2022) for the inductive settings, where new nodes appear after the training process.
Baselines. Both NodeDup and NodeDup(L) are flexible to integrate with different GNN encoder architectures and LP decoders. For our experiments, we use GSage (Hamilton et al., 2017) encoder and the inner product decoder as the default base LP model. To comprehensively evaluate our work, we compare NodeDup against three categories of baselines. (1) Base LP models. (2) Cold-start methods: TailGNN (Liu et al., 2021) and Cold-brew (Zheng et al., 2021) primarily aim to enhance the performance on cold nodes. We also compared with Imbalance (Lin et al., 2017), viewing cold nodes as an issue of the imbalance concerning node degrees. (3) Graph data augmentation methods: Augmentation frameworks including DropEdge (Rong et al., 2019), TuneUP (Hu et al., 2022), and LAGNN (Liu et al., 2022b) typically improve the performance while introducing additional preprocessing or training time.
@@ 4.1 Experimental Settings
4.2 Performance Compared to Base GNN LP Models
Isolated and Low-degree Nodes. We compare our methods with base GNN LP models that consist of a GNN encoder in conjunction with an inner product decoder and are trained with a supervised loss. From Table 1, we observe consistent improvements for both NodeDup(L) and NodeDup over the base GSage model across all datasets, particularly in the Isolated and Low-degree node settings. Notably, in the Isolated setting, NodeDup achieves an impressive 29.6% improvement, on average, across all datasets. These findings provide clear evidence that our methods effectively address the issue of sub-optimal LP performance on cold nodes.
Warm Nodes and Overall. It is encouraging to see that NodeDup(L) consistently outperforms GSage across all the datasets in the Warm nodes and Overall settings. NodeDup also outperforms GSage in 13 out of 14 cases under both settings. These findings support the notion that our methods can effectively maintain and enhance the performance of Warm nodes.
@@ 4.3 Performance Compare to Cold-start Methods
Table 1 presents the LP performance of various cold-start baselines. For both Isolated and Low-degree nodes, we consistently observe substantial improvements of our NodeDup and NodeDup(L) methods compared to other cold-start baselines. Specifically, NodeDup and NodeDup(L) achieve 38.49% and 34.74% improvement for Isolated nodes on average across all datasets, respectively.
In addition, our methods consistently outperform cold-start baselines for Warm nodes across all the datasets, where NodeDup(L) and NodeDup achieve 6.76% and 7.95% improvements on average, respectively. This shows that our methods can successfully overcome issues with degrading performance on Warm nodes in cold-start baselines.
Table 2: Performance in inductive settings (evaluated by Hits@10). The best result is bold, and the runner-up is underlined. Our methods consistently outperform GSage.

### v2

@@ 3.2 How does Node Duplication Help Cold-start LP?
Figure 4: Ablation study of NodeDup on Physics. Both Step II and Step III, introduced in Algorithm˜1, play an important role in performance improvements of NodeDup.
Ablation Study. Figure˜4 shows an ablation study on these two designs where NodeDup w/o Step III indicates only using the augmented nodes and edges in aggregation but not supervision; NodeDup w/o Step II indicates only using the augmented edges in supervision but not aggregation. We can observe that using augmented nodes and edges either in supervision or aggregation can significantly improve the LP performance on Isolated nodes. By combining them, NodeDup results in larger improvements. Besides, NodeDup also achieves improvements on Low-degree nodes while not sacrificing the performance on Warm nodes.
@@
3.3 Further Insight: Understanding NodeDup through Self-distillation
@@
As introduced in Section˜3.2, the effectiveness of NodeDup arises from leveraging diverse perspectives or signals obtained during both the aggregation and supervision steps. Allen-Zhu & Li (2020) showed that the success of self-distillation, similar to our method, contributes to ensemble learning by providing models with different perspectives on the knowledge. Building on this insight, we show an interesting interpretation of NodeDup, positioning it as a simplified and enhanced adaptation of self-distillation for LP tasks for cold nodes, as illustrated in Figure˜5, in which we draw a connection between self-distillation and NodeDup.
@@
Figure 6: Performance with different training strategies introduced in Figure˜5 on Citeseer. NodeDup achieves better performance across all settings.
@@
In self-distillation, a teacher GNN is first trained to learn the node representations $\mathbf{H}^{t}$ from original features $\mathbf{X}$ through supervised training for LP tasks on the original graph. We denote the original graph as $G^{o}$, and we denote the graph, where we replace the node features in $G^{o}$ with $\mathbf{H}^{t}$, as $G^{t}$ in Figure˜5. The student GNN is then initialized with random parameters and trained with the sum of two loss functions: $\mathcal{L}_{SD}=\mathcal{L}_{SP}+\mathcal{L}_{KD}$, where $\mathcal{L}_{SP}$ denotes the supervised training loss with $G^{o}$ and $\mathcal{L}_{KD}$ denotes the knowledge distillation loss with $G^{t}$. Figure˜6 shows that self-distillation outperforms the teacher GNN across all settings.
@@
The effect of $\mathcal{L}_{KD}$ is similar to that of creating an additional link connecting nodes in $G^{o}$ to their corresponding nodes in $G^{t}$ when optimizing with $\mathcal{L}_{SP}$. This is illustrated by the red dashed line in Figure˜5. For better clarity, we show the similarities between these two when we use the inner product as the decoder for LP with the following example. Given a node $v$ with normalized teacher embedding $\bm{h}_{v}^{t}$ and normalized student embedding $\bm{h}_{v}$, the additional loss term that would be added for distillation with cosine similarity is $\mathcal{L}_{KD}=-\frac{1}{N}\sum_{v\in\mathcal{V}}\bm{h}_{v}\cdot\bm{h}_{v}^{t}$. On the other hand, for the dashed line edges in Figure˜5, we add an edge between the node $v$ and its corresponding node $v^{\prime}$ in $G^{t}$ with embedding $\bm{h}_{v^{\prime}}^{t}$. When trained with an inner product decoder and binary cross-entropy loss, it results in the following: $\mathcal{L}_{SP}=-\frac{1}{N}\sum y_{vv^{\prime}}\log(\bm{h}_{v}\cdot\bm{h}_{v^{\prime}}^{t})+(1-y_{vv^{\prime}})\log(1-\bm{h}_{v}\cdot\bm{h}_{v^{\prime}}^{t})$. Since we always add the edge $(v,v^{\prime})$, we know $y_{vv^{\prime}}=1$, and can simplify the loss as follows: $\mathcal{L}_{SP}=-\frac{1}{N}\sum\log(\bm{h}_{v}\cdot\bm{h}_{v^{\prime}}^{t})$. Here, we can observe that $\mathcal{L}_{KD}$ and $\mathcal{L}_{SP}$ are positively correlated as $\log(\cdot)$ is a monotonically increasing function.
@@
To further improve this step and mitigate potential noise in $G^{t}$, we explore a whole graph duplication technique, where $G^{t}$ is replaced with an exact duplicate of $G^{o}$ to train the student GNN. The results in Figure˜6 demonstrate significant performance enhancement achieved by whole graph duplication compared to self-distillation. NodeDup is a lightweight variation of the whole graph duplication technique, which focuses on duplicating only the cold nodes and adding edges connecting them to their duplicates. From the results, it is evident that NodeDup consistently outperforms the teacher GNN and self-distillation in all scenarios. Additionally, NodeDup exhibits superior performance on isolated nodes and is much more efficient compared to the whole graph duplication approach.
@@ 3.4 NodeDup(L): An Efficient Variant of NodeDup
Figure 7: Aggregation and supervision mechanisms of NodeDup(L).
Inspired by the above analysis, we further introduce a lightweight variant of NodeDup for better efficiency, NodeDup(L). To provide above-described “multi-view” information as well as the supervision signals for cold nodes, NodeDup(L) simply add additional self-loop edges for the cold nodes into the edge set $\mathcal{E}$, that is, =$\mathcal{E}^{\prime}=\mathcal{E}\cup\{e_{vv}:\forall v\in\mathcal{V}_{cold}\}$. During aggregation, NodeDup(L) intentionally incorporates the self-representation $\bm{h}_{v}^{(l)}$ into the aggregated neighbors’ representation $\bm{m}^{(l)}_{v}$ through these additional edges. This allows the weight matrix $\text{W}_{2}$ to provide an extra “view” of $\bm{h}_{v}^{(l)}$ when updating $\bm{h}_{v}^{(l+1)}$. For supervision, the added edges also serve as positive training samples for cold nodes. As demonstrated in Figure˜7, NodeDup(L) preserves the two essential designs of NodeDup while avoiding the addition of extra nodes, which further saves time and space complexity. Moreover, NodeDup differs from NodeDup(L) since each duplicated node in NodeDup will provide another view for itself because of dropout layers, which leads to different performance as shown in Section˜4.2.
@@ 4 Experiments
4.1 Experimental Settings
Datasets and Evaluation Settings. We conduct experiments on 7 benchmark datasets: Cora, Citeseer, CS, Physics, Computers, Photos and IGB-100K, with their details specified in Appendix˜B. We randomly split edges into training, validation, and testing sets. We allocated 10% for validation and 40% for testing in Computers and Photos, 5%/10% for testing in IGB-100K, and 10%/20% in other datasets. We follow the standard evaluation metrics used in the Open Graph Benchmark (Hu et al., 2020) for LP, in which we rank missing references higher than 500 negative reference candidates for each node. The negative references are randomly sampled from nodes not connected to the source node. We use Hits@10 as the main evaluation metric (Han et al., 2022). We follow Guo et al. (2022) and Shiao et al. (2022) for the inductive settings, where new nodes appear after the training process. Additionally, results for large-scale datasets and heterophilic graphs are presented in Section˜C.3 and Section˜C.4.
Baselines. Both NodeDup and NodeDup(L) are flexible to integrate with different GNN encoder architectures and LP decoders. For our experiments, we use GSage (Hamilton et al., 2017) encoder and the inner product decoder as the default base LP model. To comprehensively evaluate our work, we compare NodeDup against three categories of baselines. (1) Base LP models. (2) Cold-start methods: TailGNN (Liu et al., 2021) and Cold-brew (Zheng et al., 2021) primarily aim to enhance the performance on cold nodes. We also compared with Imbalance (Lin et al., 2017), viewing cold nodes as an issue of the imbalance concerning node degrees. (3) Graph data augmentation methods: Augmentation frameworks including DropEdge (Rong et al., 2019), TuneUP (Hu et al., 2022), and LAGNN (Liu et al., 2022b) typically improve the performance while introducing additional preprocessing or training time. Performance comparisons with heuristic methods and additional cold-start methods (e.g. Upsampling, DegFairGNN (Liu et al., 2023), SAILOR (Liao et al., 2023) and GRADE (Wang et al., 2022)) are in Section˜C.6 and Section˜C.7.
Table 1: Performance compared with base GNN and baselines for cold-start methods. The best result is bold, and the runner-up is underlined. NodeDup and NodeDup(L) outperform GSage and cold-start baselines almost all the cases.
@@ 4.1 Experimental Settings
4.2 Performance Compared to Base GNN LP Models
Isolated and Low-degree Nodes. We compare our methods with base GNN LP models that consist of a GNN encoder in conjunction with an inner product decoder and are trained with a supervised loss. From Table˜1, we observe consistent improvements for both NodeDup(L) and NodeDup over the base GSage model across all datasets, particularly in the Isolated and Low-degree node settings. Notably, in the Isolated setting, NodeDup achieves an impressive 29.6% improvement, on average, across all datasets. These findings provide clear evidence that our methods effectively address the issue of sub-optimal LP performance on cold nodes.
Warm Nodes and Overall. It is encouraging to see that NodeDup(L) consistently outperforms GSage across all the datasets in the Warm nodes and Overall settings. NodeDup also outperforms GSage in 13 out of 14 cases under both settings. These findings support the notion that our methods can effectively maintain and enhance the performance of Warm nodes. Why? The superior performance on Warm nodes is directly tied to our focus on LP tasks, where we evaluate node pair outcomes. Given the substantial number of Warm-Cold node pairs under prediction, these outcomes contribute to the overall performance metrics for both Warm node prediction. Better learning of Cold nodes thus boosts Cold-Warm node pairs link prediction performance, which subsequently elevates the prediction accuracy for Warm nodes. A more detailed experimental analysis is provided in Section˜4.8.
@@ 4.3 Performance Compare to Cold-start Methods
Table˜1 presents the LP performance of various cold-start baselines. For both Isolated and Low-degree nodes, we consistently observe substantial improvements of our NodeDup and NodeDup(L) methods compared to other cold-start baselines. Specifically, NodeDup and NodeDup(L) achieve 38.49% and 34.74% improvement for Isolated nodes on average across all datasets, respectively.
In addition, our methods consistently outperform cold-start baselines for Warm nodes across all the datasets, where NodeDup(L) and NodeDup achieve 6.76% and 7.95% improvements on average, respectively. This shows that our methods can successfully overcome issues with degrading performance on Warm nodes in cold-start baselines. Further analyses with other cold-start methods and efficiency comparisons can be found in Section˜C.7 and Section˜C.8.
4.4 Performance Compared to Augmentation Methods
Effectiveness Comparison. Since NodeDup and NodeDup(L) use graph data augmentation techniques, we compare them to other data augmentation baselines. The performance and time consumption results for Citeseer, Physics, and IGB-100K are shown in Figure˜8, with additional datasets in Section˜C.9 due to space constraints. From Figure˜8, NodeDup consistently outperforms all the graph augmentation baselines for Isolated and Low-degree nodes across all three datasets, while NodeDup(L) outperforms baselines in 17/18 cases for Isolated and Low-degree nodes. Both NodeDup and NodeDup(L) also perform on par or above baselines for Warm nodes.
Efficiency Comparison. Augmentation methods often come with the trade-off of adding additional run time before or during model training. For example, LAGNN (Liu et al., 2022b) requires extra preprocessing time to train the generative model prior to GNN training. It also takes additional time to generate extra features for each node during training. Although Dropedge (Rong et al., 2019) and TuneUP (Hu et al., 2022) are free of preprocessing, they require additional time to drop edges in each training epoch compared to base GNN training. Furthermore, the two-stage training employed by TuneUP doubles the training time compared to one-stage training methods. For NodeDup methods, duplicating nodes and adding edges is remarkably swift and consumes significantly less preprocessing time than other augmentation methods. As an example, NodeDup(L) and NodeDup are 977.0$\times$ and 488.5$\times$ faster than LAGNN in preprocessing Citeseer, respectively. We also observe that NodeDup(L) has the least training time among all augmentation methods and datasets, while NodeDup also requires less training time in 8/9 cases. Additionally, NodeDup(L) achieves significant efficiency benefits compared to NodeDup in Figure˜8, especially when the number of nodes in the graph increases substantially. Taking the IGB-100K dataset as an example, NodeDup(L) is 1.3$\times$ faster than NodeDup for the entire training process.
Table 2: Performance in inductive settings. The best result is bold, and the runner-up is underlined. Our methods consistently outperform GSage.